# Closed-loop Long-horizon Robotic Planning via Equilibrium Sequence Modeling

### v5

@@ 2 Related Work
LLMs for Planning. Scaling up inference computation to improve LLMs’ performance on planning and reasoning tasks has received increasing attention (Brown et al., 2024; Snell et al., 2025; Wu et al., 2025; Jaech et al., 2024; Guo et al., 2025). Precedent techniques involving chain-of-thought (Wei et al., 2022; Zelikman et al., 2022, 2024), repeated sampling (Wang et al., 2023; Brown et al., 2024) and tree search (Yao et al., 2023a; Zhao et al., 2023) showed preliminary results with handcraft system 2. Alternatively, a method called self-refinement (Welleck et al., 2023; Shinn et al., 2023; Kim et al., 2023b; Madaan et al., 2023) suggests recursively refining the existing LLM output in an autonomous manner, but it relies heavily on prompting or sophisticated reinforcement learning. To fully exploit its potential, we propose an end-to-end optimization method for self-refinement via deep equilibrium models.
@@ 3 Method
We study the problem of robot task planning which aims to decompose a high-level task description into long-horizon mid-level action sequences (Kaelbling & Lozano-Pérez, 2011). In the following, we first discuss the limitations of LLM planners in self-refinement (Section 3.1) and address them with a novel equilibrium sequence modeling scheme (Section 3.2). This framework is adapted to closed-loop feedback with efficient designs in Section 3.3. Finally, practical implementations are presented in Section 3.4.
@@ 3.2 Self-Refinement as An Equilibrium Model
(a) Inference: solving equilibrium via fixed-point iteration
(b) Training
@@ 3.3 Equilibrium Models with Feedback
Figure 3: Illustration of our proposed framework. It incorporates (1) a memory containing all equilibrium experiences during inference, (2) a self-refining planner trained on equilibrium plans alongside the ground truth, and (3) a world model trained on experiences to simulate environmental feedback. Three modules work synergistically for closed-loop planning.
@@
3.4 Practical Implementation
This section discusses the implementation of the proposed equilibrium planner. To enable effective training while interacting with the environment, the following two modules are carefully devised to complement the planner: an experience memory that caches all equilibrium plans and their feedback during equilibrium solving, and a world model to estimate the feedback in the absence of environmental interactions. The complete planning framework is illustrated in Figure 3.
@@ 3.4 Practical Implementation
4 Experiments
4.1 Experimental Settings
@@ 5 Conclusion
Acknowledgements. We thank Xingjian Bai and Liyuan Wang for their helpful discussions.
@@ Appendix C Additional Results
Table 5: Number of iterations for LLMs to reach a fixed point.
Mean Std Min Max
Original Llama 6.70 2.12 3.22 14.98
Supervised finetuned Llama 2.46 0.88 2.02 3.80
Ours 3.02 1.61 2.32 4.07
Fixed-point convergence. Table 5 shows the number of iterations for an LLM to reach its fixed point, aggregated over 10 initial plan for each of 60 random tasks. All LLMs considered, including the original Llama, the supervised finetuned Llama, and our model, can converge to a fixed point within a few iterations, regardless of the initial sequence. This is partly due to our greedy sampling strategy (described in Section B.3) that reduces the LLMs’ randomness, after which they tend to repeat themselves (please see Figures 4(a), 10(a) and 11(a)) and converge easily.
@@ Appendix D Limitations
In a broader sense, our method may be limited in generalizing to new domains because it requires the ground truth and environmental feedback during training. These procedures with the equilibrium solving process results in lower training efficiency. Also, the current formulation only considers the explicit output plan without implicit reasoning steps. While it is possible to synergize planning and reasoning with their combined advances, we leave this to future work due to feasibility constraints for small-scale experiments with large reasoning LLMs (Jaech et al., 2024; Guo et al., 2025). Furthermore, our model has only text input and no visual input, which limits its applicability in the real world. This can be resolved by introducing video-based planners (Du et al., 2024) and world models (Yang et al., 2024; Brooks et al., 2024).

### v6

@@ 2 Related Work
LLMs for Planning. LLMs demonstrate outstanding capabilities in robotic planning (Silver et al., 2024; Zhang et al., 2023; Nayak et al., 2024; Wang et al., 2024). Scaling up inference computation to improve LLMs’ performance on planning and reasoning tasks has received increasing attention (Brown et al., 2024; Snell et al., 2025; Wu et al., 2025; Jaech et al., 2024; Guo et al., 2025). Precedent techniques involving chain-of-thought (Wei et al., 2022; Zelikman et al., 2022, 2024), repeated sampling (Wang et al., 2023; Brown et al., 2024) and tree search (Yao et al., 2023a; Zhao et al., 2023) showed preliminary results with handcraft system 2. Alternatively, a method called self-refinement (Welleck et al., 2023; Shinn et al., 2023; Kim et al., 2023b; Madaan et al., 2023) suggests recursively refining the existing LLM output in an autonomous manner, but it relies heavily on prompting or sophisticated reinforcement learning. To fully exploit its potential, we propose an end-to-end optimization method for self-refinement via deep equilibrium models.
@@ 3 Method
We study the problem of robot task planning that aims to decompose a high-level task description into long-horizon mid-level action sequences (Kaelbling & Lozano-Pérez, 2011). In the following, we first introduce the closed-loop robotic planning problems (Section 3.1). Then we discuss the limitations of LLM planners in self-refinement (Section 3.2) and address them with a novel equilibrium sequence modeling scheme (Section 3.3). This framework is adapted to closed-loop feedback with efficient designs in Section 3.4. Finally, practical implementations are presented in Section 3.5.
3.1 Problem statement
We assume that the agent runs in a fully observed environment with coarse feedback. Given a high-level task described in natural language instruction $I_{h}$ and the corresponding environment information $Env$, the agent is required to decompose $I_{h}$ into a sequence of low-level actions $\{a_{1},a_{2}...a_{n}\}$ as an action plan to accomplish its sub-goals $\{g_{1},g_{2},...g_{m}\}$, where $a_{n}$ are semantic actions from the given action space and $g_{m}$ are goal-oriented conditions that are invisible to the agent. In terms of closed-loop task planning, we allow the agent to interact with the environment many times and receive feedback for adjustment.
@@ 3.3 Self-Refinement as An Equilibrium Model
(a) Inference: solving equilibrium via fixed-point iteration
(b) Training
@@
Figure 3: Illustration of our proposed framework. It incorporates (1) a memory containing all equilibrium experiences during inference, (2) a self-refining planner trained on equilibrium plans alongside the ground truth, and (3) a world model trained on experiences to simulate environmental feedback. Three modules work synergistically for closed-loop planning. The planner interacts with the environment to output $\text{plan}_{\text{t}}$ with equilibrium sequence modeling and the equilibrium plans with feedback are stored in the equilibrium memory. During training, The planner is trained on these equilibrium plans together with the ground truth while the world model is trained on the past experiences stored in the equilibrium memory.
3.4 Equilibrium Models with Feedback
@@
3.5 Practical Implementation
This section discusses the implementation of the proposed equilibrium planner. To enable effective training while interacting with the environment, the following two modules are carefully devised to complement the planner: an experience memory that caches all equilibrium plans and their feedback during equilibrium solving, and a world model to estimate the feedback in the absence of environmental interactions. The complete planning framework is illustrated in Figure 3 and the specific implementation details are explained below.
@@ 4.4 Ablation Study
Fixed-point convergence. Table 4 shows the number of iterations for an LLM to reach its fixed point, aggregated over 10 initial plan for each of 60 random tasks. All LLMs considered, including the original Llama and Qwen, the supervised finetuned Llama, and our model, can converge to a fixed point within a few iterations, regardless of the initial sequence. This is partly due to our greedy sampling strategy (described in Section B.3) that reduces the LLMs’ randomness, after which they tend to repeat themselves (please see Figures 4(a), 10(a) and 11(a)) and converge easily.
Table 4: Convergence analysis. They summarize the number of iterations for LLMs to reach a fixed point over 10 runs on 60 random tasks with the equilibrium solving prompts. Each of them can reach convergence in a few steps.
Mean Std Min Max
Original Llama-3-8B 6.70 2.12 3.22 14.98
Original Qwen-2.5-0.5B 4.80 2.13 2.69 9.43
Original Qwen-2.5-7B 7.68 5.03 2.46 16.78
Supervised Finetuned Llama 2.46 0.88 2.02 3.80
Ours 3.02 1.61 2.32 4.07
Robustness to noisy feedback. Table 5 shows the robustness of our model when the environment is disturbed by noise. The case where the model only receives environmental feedback is considered. During inference, we randomly replace some of the feedback with incorrect feedback, such as incorrect feedback types or incorrect corresponding objects and actions. As shown in the results, our model exhibits stable performance under small amounts of noise ($\leq$10%), demonstrating its robustness in a disturbed environment.
@@ 5 Conclusion
Table 5: Robustness to disturbed environment. During inference, random feedback is injected into the environmental feedback according to the noise ratio. Our model remains stable, demonstrating robustness to noisy feedback.
Both Novel Novel Scene Novel Task
Noise ratio SR GCR SR GCR SR GCR
0% 51.61 75.13 75.79 85.79 56.62 75.53
5% 51.61 74.30 75.79 85.40 54.27 73.89
10% 53.23 73.43 74.74 85.84 53.85 73.09
20% 50.00 73.10 73.68 83.22 54.49 71.80
@@
Acknowledgements. The work is supported by an internal grant of Peking University (2024JK28), a grant from China Tower Corporation Limited. We thank Xingjian Bai and Liyuan Wang for their helpful discussions.
@@ Appendix C Additional Results
Table 7: Zero-shot evaluation on ALFRED. Our model demonstrates better generalization without retraining.
Task classification acc. Action-object recall
SFT Llama 11% 0.50%
Ours 54% 27.08%
Generalization to environment without feedback. we test our pre-trained VirtualHome planner on ALFRED benchmark with updated prompts. For evaluation, we consider two planning metrics: task classification accuracy (across 7 task types in ALFRED) and recall of ground-truth action-object pairs in the predicted plan. As shown in  Table 7, our planner generalizes significantly better than the supervised trained planner.
@@ Appendix D Limitations
In a broader sense, our method may be limited in generalizing to new domains because it requires the ground truth and environmental feedback during training. These procedures with the equilibrium solving process results in lower training efficiency. Also, the current formulation only considers the explicit output plan without implicit reasoning steps. While it is possible to synergize planning and reasoning with their combined advances, we leave this to future work due to feasibility constraints for small-scale experiments with large reasoning LLMs (Jaech et al., 2024; Guo et al., 2025). Furthermore, our model has only text input and no visual input, which limits its applicability in the real world. This can be resolved by introducing video-based planners (Du et al., 2024), world models (Yang et al., 2024a; Brooks et al., 2024) and vision-language models(Yang et al., 2024b; Athalye et al., 2024).